OpenStudy (anonymous):

What is the vertex form of the equation? y = x2 + 4x – 3

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

x^2 +4x-3 =x^2 +4x +4 -4 -3 =(x+2)^2-4-3 =(x+2)^2 -7 so the vertex is (-2,-7)